OpenStudy (goformit100):

A student multiplied 7236 by 65 instead of multiplying by 56. How many his answer was greater than the correct answer ? ( Hint: You do not have to do both multiplications ).

hero (hero):

It was greater by 65124

hero (hero):

I didn't do both multiplications. The calculator did.

hero (hero):

But if there's something you wish to show us, just go ahead and show it.

OpenStudy (goformit100):

What is the actual conception behind it ?

hero (hero):

I figured it out the conventional way. (7236)(65) - (7236)(56) = (7236)(65 - 56) = (7236)(9) = 65124